If two trucks leave a warehouse at the same time. One travels south at 61 miles per hour. The other travels north at 45 miles per hour how many hours will the trucks be at 795 miles apart?
Answer by rfer(16322)

trucks are moving apart at 61+45=106 mph
t=D/r
t=795/106
t=7.5 hrs or 7 hrs 30 min